OpenStudy (anonymous):

Fungai steps into the supermarket to buy a candle. A candle costs 29 cents. She has a $2 note, and the shopkeeper has 5 cent and 6 cent coins only in his till. If Fungai is to get her change in these denominations, what is the largest product she can obtain after multiplying the number of 5 cent coins by the number of 6 cent coins in her change? The solution starts with the equation 6x + 5y = 171. Where did they get 171

OpenStudy (anonymous):

171 +29 = 200

OpenStudy (anonymous):

6/100 = .06 cents 5/100 = .05 cents 200/100 = $2.00

OpenStudy (anonymous):

.06x + .05y = 1.71 times 100 to vacate the decimal points

OpenStudy (anonymous):

If she buys just one candle, then remaining is 200 cents - 29 cents = 171 cents

OpenStudy (anonymous):

her change (1.71) consists of how many (.06) and how many (.05)?

OpenStudy (anonymous):

Do u hv the answer??? Is it 15 coins of 5 cents and 16 coins of 6 cents their product is 15 x 16 =240 ??????

OpenStudy (anonymous):

@iamignorant Is my suggested answer right???

OpenStudy (anonymous):

Yes you are right Harkirat

OpenStudy (amistre64):

6s + 5f = 171 M = sf ; f= (-6s +171)/5 M = s[(-6s +171)/5] M = (-6s^2 +171s)/5 M' = (-12s +171)/5 M' = 0 when s = 171/12 = 14.25; so its close to that when six cent = 14 or 15 looks to be the options; and Hark decided that fate :)

OpenStudy (anonymous):

My logic goes as follows: one 5 cent coin and one 6 cent coin together make 11 cents If we divided 171 by 11, we get Quotient as 15 and remainder is 6 This means that if she gets 15 coins of 5 cents and 15 coins of 6 cents then remainder is 6 cents So we can say that if she has 15 coins of 5 cents it is = 75 cents and 15 + 1 coins of 6 cents ie 16 coins of 6 cents it is = 96 cents 75 + 96 = 171 so product of number of coins is 15 x 16 = 240 Hope it explains how I arrived at the solution......!!!